OpenStudy (anonymous):

How do you factor: 6x^2 -13x + 2

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

Looking for a method rather than a solution..

OpenStudy (anonymous):

Solution is: (x-2) (6x-1)

OpenStudy (anonymous):

Now all I need is a method to get to that solution.

OpenStudy (anonymous):

you multiply the coefficient of x^2 the constant here it is 6*2=12 then the coefficient of x =-13 so all you need to get is a pair of numbers whose sum =-13 and whose product = 12 so in your case it is -12 and -1 so we will rewrite the expression as 6x^2 -12x-x+2 =6x(x-2)-1(x-2) take (x-2) common it will be = (6x-1)(x-2) hope that is what you are looking for :)

OpenStudy (anonymous):

\[6x ^{2}-13x+2\]

OpenStudy (anonymous):

(6x-1)(x-2)

OpenStudy (anonymous):

@ Iferah. Thats exactly what I was looking for! Thanks guys!